A column of air 26cm long is trapped by a Mercury thread 5 cm.whenthe tube is inverted the air column now is 28cm. What is the atmospheric pressure
1
Expert's answer
2020-02-24T10:10:12-0500

Let the atmospheric pressure be P and the pressure of the trapped gas be p

when tube is in vertical direction then

"p_1=P+5"

when tube is inverted

"p_2=P-5"

As here temperature and number of moles are constant so we can apply boyle's law

"pV=constant\newline
p_iV_i=p_fV_f"

"(P+5)*A*26=(P-5)*A*28"

"26P+130=28P-140"

"P=135cm \ of \ hg"